### v1

@@ 2.2.2 DeepSets
$T\left(\mathcal{X};\zeta_{\Lambda}\right)=a\left(\left\{\Lambda\left(\widetilde{X}_{i};\zeta_{\Lambda}\right)\right\}_{i=1}^{m}\right),$
@@
In this framework, each data point $\widetilde{X}_{i}$ is first processed through a neural network known as the summary network ($\Lambda$), which maps the raw data into a learned feature space. These transformed features are then aggregated using a permutation-invariant function $a(\cdot)$. Some examples of permutation-invariant functions include the arithmetic mean, quantiles, product and minimum or maximum. The aggregated result is then passed through a second neural network, the inference network ($\phi$), which produces the final parameter estimates.
@@ 3.1 Experimental Setup
Before proceeding, there are various design choices in the NBE setup that must be chosen, such as the number of training samples, neural network architecture, loss function, etc. Our simulation experiments are set up to guide the selection of these components. Full details of these simulation studies are provided in Appendix 6.1.
Combining the recommendations of Sainsbury-Dale et al. [2023] and our simulation studies (see Appendix 6.1), we fix $K=10,000$ parameter samples and $J=10$ datasets per parameter, yielding a training set of $100,000$ simulations (see Equation (2)). This setup provides a balance between statistical efficiency and computational cost. The input dimension $n_{t}$, the number of observations passed to the network, can be application dependent. As our applied interest is in one-minute cryptocurrency returns over individual days, we use an input size of $n_{t}=1440$, i.e., a day worth of data. Although this is easily changed for other applications, it does involve retraining the network.
One advantage of the NBE framework is its flexibility in neural network design. We use fully connected dense neural networks (DNNs) for both the summary network $\Lambda$ and the inference network $\phi$. Each network consists of 3 hidden layers with 32 neurons per layer. Although we explored a variety of alternative architectures to support this choice, it is worth noting the performance of NBE appears generally robust to moderate changes in network design, as observed in our simulation studies.
Based on the results from our tuning experiments, we selected the mean as the aggregation function, Leaky ReLU as the activation function, and the mean squared logarithmic error (MSLE) as the loss function. These choices provided the best overall performance across the different parameters considered. Finally, we assume that the model parameters are independent and place uninformative priors on each, using uniform distributions over fixed intervals.
These choices strike a balance between accuracy and computational efficiency, enabling us to focus on evaluating the estimator’s performance across different Lévy process models. The remainder of this section describes the parameter settings, simulation structure, and comparative benchmarks used in our evaluation.
@@ 3.2 Standard Lévy Models
To explore performance of the proposed approach, we compare our method against two established estimation approaches: the least squares estimator used in [Berry and Kleiber, 2023] and the MELE method. For both, we used the same grid of characteristic function frequencies.
@@
We evaluate the performance of the proposed estimator on three benchmark models: the compound Poisson process, the variance gamma process, and the Merton jump-diffusion model. See Section 6.1 for a detailed description of the prior distributions for each model. For each model, we train the neural network estimator and assess its performance on an independent test set of $1000$ datasets in which the goal is to estimate the model parameters based on simulated test data.
In these analysis, as well as in the simulation studies, we implement the NBE approach using the NeuralEstimators package, which is written in Julia but also provides an R interface [Sainsbury-Dale, 2024]. All experiments were conducted on an Apple M3 MacBook Pro with a 10-core CPU, 16-core GPU, 16-core Neural Engine, and 18 GB of unified memory.
@@
Tables 1, 2, and 3 contain the average root mean squared error (RMSE) for each parameter across the $1000$ test datasets, along with the total estimation time required by each method. Across all models and parameters, NBE consistently outperforms competing methods in terms of estimation accuracy, sometimes substantially. It achieves the lowest RMSE for every parameter in every model considered. Additionally, NBE demonstrates a significant advantage in computational efficiency. While traditional methods such as least squares and MELE require several hundred seconds – and in some cases over a one hundred thousand seconds – to process the full assessment set, NBE achieves comparable or better accuracy in just 3 to 6 seconds per model. This dramatic speed-up highlights the practical benefits of amortized inference: once trained, the neural estimator delivers high-quality parameter estimates with minimal computational cost.
@@
It is important to note that the estimation times reported in Tables 13 refer exclusively to inference time. That is, the time required to produce parameter estimates once the neural estimator has already been trained. This is a key advantage of our approach: once training is complete, inference can be performed extremely quickly. This characteristic is known as amortized learning—a framework where computational cost is concentrated upfront during a training phase.
@@
While the initial training phase can be computationally intensive, it only needs to be performed once per model specification and prior distribution. In our own experiments, training time varied depending on the complexity of the model and the choices of hyperparameters. Across all simulation settings, we observed training durations ranging from approximately 2 hours for simpler models to as much as 55 hours for more complex configurations, such as the level-2 DVG process with large $K$ and $n_{t}$. On average, training required around 30 hours. However, these training costs are offset by the significant gains in speed and scalability at inference time, which is particularly beneficial for high-frequency financial data.
Compound Poisson process Est Time[s] RMSE ($\lambda$) RMSE ($\mu$) RMSE ($\sigma^{2}$)
LSQ 290 0.044 0.31 0.14
MELE 62,110 0.40 0.38 0.10
NBE 3 0.040 0.021 0.019
@@
Table 1: Average RMSE for the jump intensity ($\lambda$), mean ($\mu$) and variance ($\sigma^{2}$) of each jump of a compound Poisson process. The table also includes the time required to estimate the parameters for the 1,000 simulated datasets. The results were obtained using $nt=500$, $K=5000$ and $J=10$.
Merton model Est Time[s] RMSE ($\mu$) RMSE ($\sigma^{2}$) RMSE ($\lambda$) RMSE ($\mu_{j}$) RMSE ($\sigma^{2}_{j}$)
LSQ 1638 0.18 0.20 0.42 0.36 0.34
MELE 169,230 0.66 0.46 0.54 1.11 0.79
NBE 4 0.10 0.11 0.25 0.19 0.24
@@
Table 2: Average RMSE for the mean ($\mu$) and variance ($\sigma^{2}$) of the diffusion component, along with the jump intensity ($\lambda$), mean ($\mu_{j}$) and variance ($\sigma_{j}^{2}$) of the jump component of a Merton model. The table also includes the time required to estimate the parameters for the 1,000 simulated datasets. The results were obtained using $nt=1000$, $K=5000$ and $J=10$.
Variance gamma process Est Time[s] RMSE ($\gamma$) RMSE ($\sigma^{2}$) RMSE ($\alpha$)
LSQ 100 0.093 0.28 0.22
MELE 67,777 0.74 0.99 1.09
NBE 6 0.067 0.12 0.16
@@
Table 3: Average RMSE for the drift ($\gamma$), variance ($\sigma^{2}$) and variance ($\alpha$) of the subordinator for a variance gamma process. The table also includes the time required to estimate the parameters for the 1,000 simulated datasets. The results were obtained using $nt=1000$, $K=5000$ and $J=10$.
@@ 3.3 Deep Variance Gamma
We test our framework on a more difficult estimation task involving multiply-subordinated processes; in particular we consider a two-level DVG model. As outlined in Section 2.1.2, the two-level DVG parameters are the variance parameter $\sigma^{2}$ and the subordination parameters $\alpha_{1},\alpha_{2}$, corresponding to each level of gamma subordination. To this end, we trained the neural estimator and evaluated its performance on 1000 datasets, each generated from parameter values sampled from the prior distribution.
We selected prior distributions to be reasonably uninformative, reflecting limited prior knowledge while remaining consistent with ranges typically used in empirical studies of financial Lévy models and allowing for sufficient flexibility in the estimation. For instance, variance parameters in Lévy models are often on the order of 1 or below, and subordination parameters typically fall within moderate positive ranges [Berry and Kleiber, 2023]. Accordingly, we chose uniform priors: $\sigma^{2}\sim U(10^{-6},3)$ and $\alpha_{1},\alpha_{2}\sim U(10^{-6},25)$, allowing the model to explore a broad but plausible range of parameter values.
Berry and Kleiber [2023] showed that potential model identifiability concerns may arise in the DVG, with certain parameter combinations exhibiting very similar implied characteristic functions, except for at either very small, or very large, frequencies. Thus, to assess the quality of our estimates, we introduce an additional metric that captures functional differences between the estimated and true models. Given the true parameter vector $\theta=\left(\sigma^{2},\alpha_{1},\alpha_{2}\right)^{\rm T}$, we compute the associated theoretical characteristic function $\Phi(\omega,\theta)$ at frequency $\omega$. Likewise, for an estimate $\hat{\theta}=\left(\hat{\sigma}^{2},\hat{\alpha}_{1},\hat{\alpha}_{2}\right)^{\rm T}$, we compute the corresponding $\Phi(\tau,\hat{\theta})$. We define the $\mathbb{L}_{2}^{f}$ distance between these vectors as
@@
Interestingly, the results also suggest a strong correlation between the two subordinator parameters, which prompted further investigation. Figure 2 displays scatter plots of $\alpha_{1}$ versus $\alpha_{2}$, both for the true parameter values and the estimates. Several insights emerge from this visualization. First, estimation appears to be more difficult when either $\alpha_{1}$ or $\alpha_{2}$ is close to zero, as indicated by higher $\mathbb{L}_{2}^{f}$ values near the bottom an left edges of the plot. More notably, the estimated values show a tendency for the model to predict $\alpha_{1}$ and $\alpha_{2}$ as being similar. This alignment suggests that the neural estimator finds its best functional fit when $\alpha_{1}\approx\alpha_{2}$, potentially revealing a structural symmetry in the underlying process.

### v2

@@ 2.2.5 DeepSets Architecture for Lévy Processes
$T\left(\mathcal{X};\zeta_{\Lambda}\right)=a\left(\left\{\Lambda\left(\widetilde{X}_{i};\zeta_{\Lambda}\right)\right\}_{i=1}^{n-1}\right),$
@@
In this framework, each data point $\widetilde{X}_{i}$ is first processed through a neural network known as the summary network ($\Lambda$), which maps the raw data into a learned feature space. These transformed features are then aggregated using a permutation-invariant function $a(\cdot)$. Some examples of permutation-invariant functions include the arithmetic mean, quantiles, product and minimum or maximum. The aggregated result is then passed through a second neural network, the inference network ($\phi$), which produces the final parameter estimates. It should be noted that, under this setup, the network is set to accommodate input dimension of size $n$, and would need to be retrained for other dimensions. Sainsbury-Dale et al. [2023] suggest some adaptations to accommodate different input dimensions, but we focus on a fixed input size for the remainder of the paper.
2.3 Theoretical Guarantees for DeepSets-based Bayes Estimation
In this section, we state a new consistency result showing that a sufficiently expressive DeepSets class trained by minimizing a Monte Carlo approximation to the Bayes risk recovers said risk.
Let $\Theta\subset\mathbb{R}^{d}$ be the parameter space and let $\mathcal{X}=(X_{1},\dots,X_{n})$ be i.i.d. replicates with distribution determined by $\theta\in\Theta$. Let $f(x\mid\theta)$ denote the data-generating density for $\mathcal{X}$ given $\theta$. Recall that a Bayes estimator (for the chosen loss) is any minimizer of the Bayes risk (2),
[…]
$R_{N}(\hat{\theta})\;=\;\frac{1}{N}\sum_{i=1}^{N}L(\theta_{i},\hat{\theta}(\mathcal{X}_{i})).$
Let $\hat{\theta}_{m,N}^{\star},\hat{\theta}_{m}^{\star}\in\mathcal{F}_{m}$ denote then empirical and population risk minimizer over $\mathcal{F}_{m}$, respectively,
$\hat{\theta}_{m,N}^{\star}\in\arg\min_{\hat{\theta_{m}}\in\mathcal{F}_{m}}R_{N}(\hat{\theta}_{m})\qquad,\qquad\hat{\theta}_{m}^{\star}\in\arg\min_{\hat{\theta}_{m}\in\mathcal{F}_{m}}R(\hat{\theta}_{m}).$
We now establish the consistency result under the following assumptions.
  1. (A1)
    Compactness, continuity and Lipschitz loss. The parameter space $\Theta$ is compact. The loss $L(\theta,a)$ is continuous on $\Theta\times\Theta$ and is uniformly Lipschitz in its second argument, i.e., there exists $0<C_{L}<\infty$ such that
    $\forall\theta\in\Theta,\ \forall a,a^{\prime}\in\Theta,\hskip 18.49988pt|L(\theta,a)-L(\theta,a^{\prime})|\leq C_{L}\|a-a^{\prime}\|.$
  2. (A2)
    Permutation invariance of Bayes rule. For i.i.d. replicates $\mathcal{X}=(X_{1},\dots,X_{n})$ the Bayes estimator $\hat{\theta}^{\star}(X)$ is a permutation–invariant function of the sample.
  3. (A3)
    DeepSets universality. The union $\bigcup_{m\geq 1}\mathcal{F}_{m}$ is dense in the space of continuous permutation-invariant functions on the sample space Zaheer et al. [2017]. This means that, for every continuous permutation-invariant $g$ on $\mathbbm{R}^{n}$ and every $\varepsilon>0$ there exists $m$ and $\hat{\theta}\in\mathcal{F}_{m}$ with $\sup_{x}\|g(x)-\hat{\theta}(x)\|<\varepsilon$.
  4. (A4)
    Uniform law of large numbers. For each fixed $m$, the class $\mathcal{F}_{m}$ satisfies,
    $\sup_{\hat{\theta}\in\mathcal{F}_{m}}\big|\hat{R}_{N}(\hat{\theta})-R(\hat{\theta})\big|\xrightarrow{\,P\,}0\qquad\text{as }N\to\infty.$
Theorem 2.
Assume (A1)–(A4). Fix a sequence $m\to\infty$ and, for each fixed $m$, let $N\to\infty$. For any sequence of empirical risk minimizers $\hat{\theta}_{m,N}^{\star}\in\mathcal{F}_{m}$, we have
$R(\hat{\theta}_{m,N}^{\star})\xrightarrow{\,P\,}R^{\star}$
for any joint sequence $\min(m,N)\to\infty$ such that $N$ increases sufficiently rapidly relative to $m$ so that Assumption (A4) holds for each class $\mathcal{F}_{m}$.
The proof of Theorem 2 can be found in the Appendix (Section 6.1 ). This theorem shows that, under a conceptually simple set of assumptions, the risk of the neural Bayes estimator converges to the Bayes risk in the DeepSets framework, as both the network capacity $m$ and the simulation budget $N$ grow.
Although this result is novel in the context of DeepSets, related progress has been made on the theoretical guarantees of neural Bayes estimators. In particular, the recent paper by Rödder et al. [2025] develops a similar risk decomposition framework to establish convergence results for a broad class of neural networks architectures. Their analysis is technically more involved and arguably more general, as it incorporates an explicit optimization error term. However, our framework has the advantage of using a general loss function in our analysis. Readers interested in sharper bounds, explicit treatment of optimization error, or a discussion of how fast $N$ must grow relative to $m$ are referred to [Rödder et al., 2025], who explore these issues in depth.
If we further strengthen our assumptions to require strong convexity of the loss function, then the Bayes minimizer is unique, and we obtain a stronger conclusion. In particular, suppose that:
  1. (A5)
    The loss function $L(\theta,a)$ is strictly convex in $a\in\Theta$.
  2. (A6)
[…]
Corollary 3.
Assume (A1)–(A6) hold, for any sequence of empirical Bayes estimators $\hat{\theta}_{m,N}^{\star}\in\mathcal{F}_{m}$, we have
$\hat{\theta}_{m,N}^{\star}\xrightarrow{\,P\,}\theta^{\star}\qquad\text{as }\min\{m,N\}\to\infty,$
provided $N\to\infty$ sufficiently fast relative to $m$.
@@ 3.1 Experimental Setup
Before proceeding, there are various design choices in the NBE setup that must be chosen, such as the number of training samples, neural network architecture, loss function, etc. Based on Theorem 1, we can, in principle, generate as many datasets as needed, so $K$ and $J$ could theoretically be chosen arbitrarily large. In practice, however, the amount of training data required depends on several factors, including the complexity of the model, the number of parameters being estimated, the chosen loss function, the neural network architecture, and even the specifics of the optimization algorithm. As a rough guideline, Sainsbury-Dale et al. [2023] suggest that values of $K$ on the order of $10^{4}$$10^{5}$ are typically necessary for the neural Bayes estimator to generalize well. Our simulation experiments are set up to guide the selection of these components. Full details of these simulation studies are provided in the Appendix (Section 6.2).
The results we present in this section are obtained using fully connected dense neural networks (DNNs) for both the summary network $\Lambda$ and the inference network $\phi$. Each network consists of 3 hidden layers with 32 neurons per layer. Although we explored a variety of alternative architectures to support this choice, it is worth noting the performance of NBE appears generally robust to moderate changes in network design, as observed in our simulation studies.
Based on the results from our tuning experiments, we selected the mean as the aggregation function, Leaky ReLU as the activation function, and the mean squared logarithmic error (MSLE) as the loss function. These choices provided the best overall performance across the different parameters considered. Finally, we assume that the model parameters are independent a priori and place uninformative priors on each, using uniform distributions over fixed intervals. These choices strike a balance between accuracy and computational efficiency, enabling us to focus on evaluating the estimator’s performance across different Lévy process models.
In the following analyses, as well as in the simulation studies, all methods were implemented in Julia to ensure comparability of computational timings. We implement the NBE approach using the NeuralEstimators package [Sainsbury-Dale, 2024]. The LSQ estimator was coded using the Optim.jl package [Mogensen and Riseth, 2018], specifically calling the optimize function. The MELE estimator was implemented using JuMP.jl [Dunning et al., 2017] with the Ipopt solver [Wächter and Biegler, 2006], where the empirical likelihood was directly maximized. Furthermore, all experiments were conducted on an Apple M3 MacBook Pro with a 10-core CPU and 18 GB of unified memory. The remainder of this section describes the parameter settings, simulation structure, and comparative benchmarks used in our evaluation.
@@ 3.2 Standard Lévy Models
To explore performance of the proposed approach, we compare our method against two established estimation approaches: the LSQ method used in [Berry and Kleiber, 2023] and the MELE method (see Section 2.2.1). For both, we used the same grid of characteristic function frequencies.
@@
We evaluate the performance of the proposed estimator on three benchmark models: the compound Poisson process, the variance gamma process, and the Merton jump-diffusion model. See Section 6.2 for a detailed description of the prior distributions for each model. For each model, we train the neural Bayes estimator and assess its performance on an independent test set of $1000$ datasets in which the goal is to estimate the model parameters based on simulated test data.
@@
Tables 1, 2, and 3 contain the average root mean squared error (RMSE), bias, and standard deviation for each parameter across the $1000$ test datasets, along with the total estimation time required by each method. Across all models and parameters, NBE consistently outperforms competing methods in terms of estimation accuracy, sometimes substantially. It achieves the lowest RMSE for every parameter in every model considered. Additionally, NBE demonstrates a significant advantage in computational efficiency. While traditional methods such as LSQ and MELE require several hundred seconds – and in some cases over a hundred thousand seconds – to process the full assessment set, NBE achieves comparable or better accuracy in just 3 to 6 seconds per model. This dramatic speed-up highlights the practical benefits of amortized inference: once trained, the neural Bayes estimator delivers high-quality parameter estimates with minimal computational cost.
@@
It is important to note that the estimation times reported in Tables 13 refer exclusively to inference time. That is, the time required to produce parameter estimates once the neural Bayes estimator has already been trained. We emphasize that NBE involves a non-trivial upfront computational cost for training the network, whereas methods such as LSQ or MELE directly optimize on each dataset without training. However, once the network has been trained, inference with NBE is essentially instantaneous, while LSQ and MELE require the solving of a new optimization problem for every data set. This is a key advantage of our approach: once training is complete, inference can be performed extremely quickly. This characteristic is known as amortized learning—a framework where computational cost is concentrated upfront during a training phase, after which model execution is fast.
@@
While the initial training phase can be computationally intensive, it only needs to be performed once per model specification and prior distribution. In our own experiments, training time varied depending on the complexity of the model and the choices of hyperparameters. Across all simulation settings, we observed training durations ranging from approximately 2 hours for simpler models to as much as 55 hours for more complex configurations, such as the level-2 DVG process with large $K$ and $n_{t}$. On average, training required around 25 hours. However, these training costs are offset by the significant gains in speed and scalability at inference time, which is particularly beneficial for high-frequency financial data.
Compound Poisson process Training [s] Est Time[s] $\lambda$ $\mu$ $\sigma^{2}$
LSQ - 833 0.058 (-0.004) [0.038] 0.23 (-0.025) [0.21] 0.13 (0.116) [0.065]
MELE - 777.7 [h]* 0.70 (-0.61) [0.38] 0.36 (-0.038) [0.19] 0.11 (-0.09) [0.067]
NBE 1276 2.8 0.051 (0.0005) [0.032] 0.028(-0.0055)[0.018] 0.022(7.6e-5) [0.014]
@@
Table 1: RMSE, (Bias) and [SD] (averaged over 1000 simulated datasets from the prior) for the jump intensity ($\lambda$), mean ($\mu$) and variance ($\sigma^{2}$) of each jump of a compound Poisson process. The table also includes the time required to estimate the parameters over all simulations. * For MELE, only 10 datasets were computed due to its high computational cost. The reported time is extrapolated from these 10 datasets. The results were obtained using $n_{t}=500$, $K=5000$ and $J=10$.
Merton model Training[s] Est Time[s] $\mu$ $\sigma^{2}$ $\lambda$ $\mu_{j}$ $\sigma^{2}_{j}$
LSQ - 695 0.17 (-0.001) [0.14] 0.19 (0.14) [0.091] 0.46 (0.13) [0.34] 0.34 (0.007) [0.25] 0.37 (-0.056) [0.25]
MELE - 1155 [h] * 0.37 (-0.21) [0.27] 0.20 (-0.12) [0.11] 1.01 (0.43) [0.39] 0.31 (-0.14) [0.083] 0.84 (-0.69) [0.53]
NBE 2085 4 0.11 (4.5e-5) [0.072] 0.11 (-0.004) [0.068] 0.25 (0.022) [0.16] 0.19 (-0.09) [0.13] 0.24(-0.015) [0.15]
@@
Table 2: RMSE, (Bias) and [SD] (averaged over 1000 simulated datasets from the prior) for the mean ($\mu$) and variance ($\sigma^{2}$) of the diffusion component, along with the jump intensity ($\lambda$), mean ($\mu_{j}$) and variance ($\sigma_{j}^{2}$) of the jump component of a Merton model. The table also includes the time required to estimate the parameters for the 1,000 simulated datasets. * For MELE, only 10 datasets were computed due to its high computational cost. The reported time is extrapolated from these 10 datasets. The results were obtained using $n_{t}=1000$, $K=5000$ and $J=10$.
Variance gamma process Training [s] Est Time[s] $\gamma$ $\sigma^{2}$ $\alpha$
LSQ - 398 0.16 (-0.005) [0.15] 0.32 (-0.083) [0.21] 1.70 (-0.59) [0.85]
MELE - 1024[h] * 0.73 (-0.025) [0.40] 1.18 (-0.99) [0.69] 1.93 (-0.98) [0.37]
NBE 672 3.5 0.048 (-0.01) [0.030] 0.090 (-0.0085) [0.062] 0.24 (0.038) [0.17]
@@
Table 3: RMSE, (Bias) and [SD] (averaged over 1000 simulated datasets from the prior) for the drift ($\gamma$), variance ($\sigma^{2}$) and variance ($\alpha$) of the subordinator for a variance gamma process. The table also includes the time required to estimate the parameters for the 1,000 simulated datasets. * For MELE, only 10 datasets were computed due to its high computational cost. The reported time is extrapolated from these 10 datasets. The results were obtained using $n_{t}=1000$, $K=5000$ and $J=10$.
@@ 3.3 Deep Variance Gamma
We test our framework on a more difficult estimation task involving multiply-subordinated processes; in particular we consider a two-level DVG model. As outlined in Section 2.1.2, the two-level DVG parameters are the variance parameter $\sigma^{2}$ and the subordination parameters $\alpha_{1},\alpha_{2}$, corresponding to each level of gamma subordination.
Combining the recommendations of Sainsbury-Dale et al. [2023] and our simulation studies (see Appendix 6.2), we fix $K=10,000$ parameter samples and $J=10$ datasets per parameter, yielding a training set of $100,000$ simulations (see Equation (3)). The input dimension $n_{t}$, the number of observations passed to the network, can be application dependent. As our applied interest is in one-minute cryptocurrency returns over individual days, we use an input size of $n_{t}=1440$, i.e., a day worth of data. Although this is easily changed for other applications, it does involve retraining the network.
In terms of the architecture used, we used the same one as in the previous section and the same aggregation, activation, and loss function. Finally, we trained the neural Bayes estimator and evaluated its performance on 1000 datasets, each generated from parameter values sampled from the prior distribution.
We selected prior distributions to be reasonably uninformative, reflecting limited prior knowledge while remaining consistent with ranges typically used in empirical studies of financial Lévy models and allowing for sufficient flexibility in the estimation, especially when applied to multiple datasets. For instance, variance parameters in Lévy models are often on the order of 1 or smaller, and subordination parameters typically fall within moderate positive ranges [Berry and Kleiber, 2023]. Accordingly, we chose uniform priors: $\sigma^{2}\sim U(10^{-6},3)$ and $\alpha_{1},\alpha_{2}\sim U(10^{-6},25)$, allowing the model to explore a broad but plausible range of parameter values. These bounds are informed by Berry and Kleiber [2023], who showed that cryptocurrency parameter estimates consistently fell within these intervals, so the truncation reflects a combination of domain expertise and empirical evidence. Despite this, we acknowledge that even flat priors can exert influence. For general discussion of this issue see Gelman et al. [2017], and for a robustness experiment in our NBE setting, we refer the reader to Section 6.2.
@@
$\|\theta-\hat{\theta}\|_{\mathbb{L}_{2}^{f}}=\displaystyle\sqrt{\displaystyle\int_{\omega}\left(\Phi(\omega,\theta)-\Phi(\omega,\hat{\theta})\right)^{2}d\omega}$ (5)
@@
It should be noted that Bayes estimators are, in general, biased and therefore are not necessarily expected to lie on the identity line in Figure 1. In principle, we could assess the role of this bias by comparing our neural Bayes estimates with the corresponding true Bayes estimates. However, computing the exact Bayes benchmarks is not feasible for the class of models considered here, which is precisely why we adopt a likelihood-free approach.
@@
Interestingly, the results also suggest a strong correlation between the two subordinator parameters, which prompted further investigation. Figure 2 displays scatter plots of $\alpha_{1}$ versus $\alpha_{2}$, both for the true parameter values and the estimates. Several insights emerge from this visualization. First, estimation appears to be more difficult when either $\alpha_{1}$ or $\alpha_{2}$ is close to zero, as indicated by higher $\mathbb{L}_{2}^{f}$ values near the bottom an left edges of the plot. More notably, the estimated values show a tendency for the model to predict $\alpha_{1}$ and $\alpha_{2}$ as being similar. This alignment suggests that the neural Bayes estimator finds its best functional fit when $\alpha_{1}\approx\alpha_{2}$, potentially revealing a structural symmetry in the underlying process.